OpenStudy (anonymous):

please HELP! Find the volume of a cylinder with a height of 5 meters and a diameter that is 4 times the measure of the height. Use 3.14 for pi, and round your answer to the nearest hundredth. 62.8 m3 251.2 m3 678.24 m3 1570 m3

OpenStudy (anonymous):

What's the formula to find the volume of a cylinder?

OpenStudy (anonymous):

V = π r 2 h

OpenStudy (anonymous):

Yup, so now just plug in the height, radius (which is half of the diameter) and the pi and solve!

OpenStudy (anonymous):

THIS IS SO HARD

OpenStudy (nurali):

V=hpir^2 d=4h r=d/2 r=4h/2 r=2h h=5 so r=10 V=5*pi*10^2 V=(5)(100)(3.14) = ?

OpenStudy (anonymous):

its 1575

OpenStudy (anonymous):

1570

OpenStudy (anonymous):

:) thanks!

OpenStudy (nurali):

Anytime.